OpenStudy (anonymous):

A rectangle has an area of 36 square feet. The width is five feet less than the length. What is the length of the rectangle?

OpenStudy (academicgurusinc):

Define the rectangle where width = W and length= L Area of rectangle = L*W=36 w= L-5 Substituted this equation for W above. That is L*W = 36 L* (L-5) = 36 L^2 - 5L = 36 L^2 - 5L - 36 = 0 Factor this: (L-9)(L+4)=0 Thefore L= 9 or -4 Since the length cannot be a negative value, L must be 9 feet. This is true because: w= L-5= 9-5= 4 L= 9 Area=L*W = =9*4 =36 We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)